# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ VI-D Sparse MIMO ISAC
Looking ahead to future 6G mobile communication networks, MIMO systems are expected to provide finer spatial resolution, not only enhancing the SE of wireless communication but also enabling more precise wireless sensing. To tackle higher power consumption and signal processing complexity of MIMO systems, sparse arrays have been proposed as an alternative configuration to achieve a large array aperture with a small number of antennas [52][162]. Sparse MIMO arrays allow for a large virtual MIMO system for sensing [163], which still enables underdetermined target parameter estimation. However, sparse MIMO also leads to a number of challenging issues. For example, since sparse antenna geometries result in antenna spacing that exceeds half wavelength, undesired grating lobes are generated. When users or targets are located within these lobes, severe communication inter-user interference and radar angle ambiguity occur. For sparse MIMO sensing, differential or sum co-arrays have been demonstrated to effectively mitigate grating lobes and prevent angular ambiguity. However, the application of these techniques to communication or ISAC systems remains unclear. Moreover, in near-field scenarios, the the spatially nonstationary phase that results from spherical wave propagation does not allow formation of a virtual array as in the far-field case [164]. Therefore, developing methods to enhance DoFs in the near-field is a critical challenge. Sparse array design itself is a challenging problem, and more work is needed to develop designs that provide an effective balance between C&S performance.
@@ VI-E NLoS Sensing and Localization in ISAC
The vision of pervasive sensing and connectivity for future 6G wireless networks relies not only on LoS signals but also on NLoS sensing and localization. Nevertheless, achieving highly accurate localization, particularly for NLoS components, still remains an open issue. The parameter ambiguity introduced by NLoS propagation presents both challenges and opportunities in ISAC. Although it is challenging to directly perform sensing and localization based on NLoS components, hybrid localization schemes (e.g., RSS-ToA [165] and AoA-ToA [166]) integrated with prior information (e.g., CSI [167] and CKM) make NLoS sensing possible. NLoS components can significantly contribute to more precise localization services and transmission coverage. Thus, IRS [70, 71] aided ISAC is regarded as a promising approach that can enhance C&S performance by creating additional NLoS paths [168].
@@ VI-F Mutualism between ISAC and CKM
VI-G Prototyping and Experimental Validation
Proof-of-concept prototype design and experimental verification are prerequisites for realizing ISAC in practical wireless networks. Numerous efforts have been devoted to addressing this issues. Early experiments [169, 170] focused on validating and analyzing the performance of using OFDM for radar sensing, which later evolved into system prototype designs for monostatic OFDM ISAC [171, 172]. In [25], the authors developed an OFDM ISAC prototype based on the 5G NR frame structure. A bistatic ISAC prototype system using an OFDM frame-based approach for sensing tasks was validated in [37]. OFDM-based simultaneous localization and mapping (SLAM) [9], synthetic aperture radar (SAR) imaging [173], and ISAC-enabled CKM prototype [32] systems have been successful implemented. In contrast, while there are many mature experimental platforms for MIMO OFDM ISAC in far-field scenarios, more prototype designs and experimental validations are needed for massive MIMO and near-field ISAC systems as well as the network-level implementations.
@@ VII Conclusion
MIMO-OFDM is the most likely wireless technology to realize ISAC in future 6G wireless networks due to its predominant use in existing wireless systems. This paper has provided a comprehensive tutorial overview on ISAC systems enabled by MIMO-OFDM. A unified approach to modeling MIMO-OFDM ISAC systems and their enabling signal processing algorithms were presented, under both far-field and near-field channel conditions. Open problems for MIMO-OFDM ISAC that deserve further investigation were also discussed. We hope that this tutorial will provide a useful guide for researchers working on ISAC, towards efficiently integrating it into future 6G wireless networks via efficient synergy with MIMO and OFDM.

### v2

@@ VI-D Sparse MIMO ISAC
Looking ahead to future 6G mobile communication networks, MIMO systems are expected to provide finer spatial resolution, not only enhancing the SE of wireless communication but also enabling more precise wireless sensing. To tackle the higher power consumption and signal processing complexity of MIMO systems, sparse arrays have been proposed as a way to achieve a large array aperture with a small number of antennas [52][181]. Sparse MIMO arrays allow for a large virtual MIMO system for sensing [182], which still enables underdetermined target parameter estimation. However, sparse MIMO also leads to a number of challenging issues. For example, since sparse antenna geometries result in antenna spacing that exceeds one half wavelength, undesired grating lobes are generated. When users or targets are located within these lobes, severe communication inter-user interference and radar angle ambiguity occur. For sparse MIMO sensing, differential or sum co-arrays have been demonstrated to effectively mitigate grating lobes and prevent angular ambiguity. However, the application of these techniques to communication or ISAC systems remains unclear. Moreover, in near-field scenarios, the spatially nonstationary phase that results from spherical wave propagation does not allow formation of a virtual array as in the far-field case [183]. Therefore, developing methods to enhance DoFs in the near-field is a critical challenge. Sparse array design itself is a challenging problem, and more work is needed to develop designs that provide an effective balance between C&S performance.
@@ VI-E NLoS Sensing and Localization in ISAC
The vision of pervasive sensing and connectivity for future 6G wireless networks relies not only on LoS signals but also on NLoS sensing and localization. Nevertheless, achieving highly accurate localization in NLoS scenarios remains an open issue. The parameter ambiguity introduced by NLoS propagation presents both challenges and opportunities in ISAC. Although it is challenging to directly perform sensing and localization based on NLoS signal components, hybrid localization schemes (e.g., RSS-ToA [184] and AoA-ToA [185]) integrated with prior information (e.g., CSI [186] and CKM) make NLoS sensing possible. NLoS components can significantly contribute to more precise localization services and transmission coverage. Thus, IRS [70, 71] aided ISAC is regarded as a promising approach that can enhance C&S performance by creating additional NLoS paths [187].
@@ VI-F Mutualism between ISAC and CKM
VI-G ISAC for 6G System-Level Requirements
Beyond signal processing, several system-level requirements for 6G networks also need to be considered for ISAC system design. For example, 6G is likely to feature ultra-dense deployment of BSs and access points (APs), requiring efficient coordination to ensure seamless connectivity and consistent sensing coverage. ISAC can facilitate such coordination by providing shared sensing information across nodes, enabling cooperative beam management and multi-static sensing [188, 189]. Many envisioned applications, such as autonomous driving and industrial automation, impose ultra-low latency requirements, necessitating efficient ISAC designs to support real-time operation [190]. By integrating sensing feedback directly into communication protocols, ISAC enables real-time perception-control loops that are difficult to achieve with conventional separated systems. Furthermore, with the anticipated deployment of massive numbers of devices in 6G, energy consumption becomes a critical bottleneck. ISAC inherently enhances efficiency by sharing spectrum, hardware, and infrastructure between communication and sensing functions, thereby avoiding redundant deployments and significantly reducing energy and cost overhead. These aspects showcase that ISAC is not merely a physical-layer technique focused on waveforms and algorithms, but also a key enabler for achieving broader 6G network objectives. Therefore, addressing these system-level requirements in a structured manner is an important direction for future ISAC research.
VI-H Security and Privacy in MIMO-OFDM ISAC
Although significant advantages in spectral efficiency and overall performance can be achieved by MIMO OFDM-aided ISAC systems, security and privacy concerns are becoming increasingly prominent [191, 192]. Communication security faces new challenges since the signal used for sensing may carry confidential information intended for communication users. Targets may thus act as eavesdroppers, threatening the confidentiality of the communication links. In an ISAC scenario, the transmitter should focus signal energy towards targets to obtain strong echoes while preventing the eavesdroppers from exploiting this signal to decode confidential information, which requires a careful trade-off [193]. Radar privacy is also a critical concern, involving not only protection of target parameter data (e.g., location, velocity) but also sensitive information related to the radar itself, such as its location and transmitted waveforms. Even when communication data is protected through encryption, eavesdroppers can still infer the AP’s location by detecting the existence of a communication link, thereby exposing the system to potential security threats. Therefore, security and privacy challenges arise in both the communication and sensing domains, which in ISAC systems are inherently coupled. This necessitates integrated protection strategies that enhance communication security while preserving sensing privacy [194].
@@ VII Conclusion
MIMO-OFDM is the most promising wireless technology to realize ISAC in future 6G wireless networks due to its predominant use in existing wireless systems. This paper has provided a comprehensive tutorial overview of ISAC systems enabled by MIMO-OFDM. A unified approach to modeling MIMO-OFDM ISAC systems and their enabling signal processing algorithms were presented, under both far-field and near-field channel conditions. Open problems for MIMO-OFDM ISAC that deserve further investigation were also discussed. We hope that this tutorial will provide a useful guide for researchers working on ISAC, towards efficiently integrating it into future 6G wireless networks via effective synergy with MIMO and OFDM.